OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 2x − 3y = 6. slope = 2/3 and y-intercept = −2 slope = −2/3 and y-intercept = −2 slope = 2/3 and y-intercept = 2 slope = −2/3 and y-intercept = 2

pooja195 (pooja195):

Start by subtracting 2x from both sides -3y=2x+6 divide all by -3 what are you left with?

OpenStudy (anonymous):

after you do all what @pooja195 wrote you should get this: -3y=6-2x then, y=6/(-3) - 2/(-3)x y=-2 + 2/3x so the line you get has slope-intercept form equation: y=2/3x - 2

pooja195 (pooja195):

@ksanka a note for future reference as a fellow helper: sometimes it's best to simply give an explanation and let the person try to find the answer themselves, otherwise they could simply glance at your comment, find the answer they were looking for, thank you, and not even read the explanation you worked so hard to type out :)